The scores of 15 students enrolled in GEC 14 class were recorded in ascending order as follows: 4, 7, 7, 9, 10, 11, 13, 15, 15, 17, 17, 17, 19, 19 and 20. After calculating the mean, median and mode, an error is discovered that one of the two 15s is really a 13. Which of the following measures of central tendency will change?

1
Expert's answer
2021-11-18T06:17:50-0500

Old data : 4, 7, 7, 9, 10, 11, 13, 15, 15, 15, 17, 17, 19, 19, 20

Mean :

Mean =13.2

Median =middle number of the obs

Median=(n+1)/2 th obs

Median=(16/2)th obs

Median=8th obs

Median =15

Mode:Highest number of frequency obs.

Mode=15 [occur highest times]

New data (corrected data)

4, 7, 7, 9, 10, 11, 13, 15, 15, 15, 17, 17, 17, 19, 20

Mean :

Median =middle number of the obs

Median=(n+1)/2 th obs

Median=(16/2)th obs

Median=8th obs

Median =15

Mode:Highest number of frequency obs.

Mode=15,17 [occur highest times]

The measures of central tendency which will change are mean and mode